# Safeguard Fine-Tuned LLMs Through Pre- and Post-Tuning Model Merging

### v1

@@ Abstract
Fine-tuning large language models (LLMs) for downstream tasks is a widely adopted approach, but it often leads to safety degradation in safety-aligned LLMs. Currently, many solutions address this issue by incorporating additional safety data, which can be impractical in many cases. In this paper, we address the question: How can we improve downstream task performance while preserving safety in LLMs without relying on additional safety data? We propose a simple and effective method that maintains the inherent safety of LLMs while enhancing their downstream task performance: merging the weights of pre- and post-fine-tuned safety-aligned models. Experimental results across various downstream tasks, models, and merging methods demonstrate that this approach effectively mitigates safety degradation while improving downstream task performance, offering a practical solution for adapting safety-aligned LLMs.
@@ 1 Introduction
The rapid advancement and increasing accessibility of Large Language Models (LLMs) necessitate a critical focus on aligning these technologies with human values, cultural norms, and trustworthiness Huang et al. (2023). To address these challenges, researchers and developers have introduced safety techniques such as preference tuning Ouyang et al. (2022); Rafailov et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024), aimed at preventing LLMs from generating harmful or inappropriate content. Consequently, an increasing number of applications leverage safety-aligned models—referred to as base models in this paper—that are further customized for downstream tasks via Supervised Fine-Tuning (SFT) Chung et al. (2024).
Figure 1: Beyond standard SFT for downstream task adaptation, our proposed method effectively mitigates safety degradation by combining the base and the fine-tuned model while simultaneously enhancing downstream task performance.
However, recent studies Yang et al. (2023); Qi et al. (2024); Zhan et al. (2024) highlight a critical challenge: fine-tuning base models can degrade their safety, even when using benign datasets. To address this issue, mainstream approaches often incorporate additional safety data during fine-tuning Qi et al. (2024); Bianchi et al. (2024). Additionally, some methods that utilize model merging techniques Hazra et al. (2024); Hsu et al. (2024) rely on auxiliary safety or harmful datasets in their complex re-alignment processes. These approaches are often constrained by the scarcity of safety data and can be resource-intensive, requiring datasets with labeled safe examples in specific domains.
Inspired by Wortsman et al. (2022b), we propose a simple yet effective method to improve downstream task performance while mitigating safety degradation. As shown in Figure 1, our approach consists of two steps: first, we fine-tune the base model on the downstream task; second, we merge the base model with the fine-tuned model. We evaluate this approach across various models, downstream tasks, and merging techniques. The experimental results demonstrate that our method enhances downstream task performance while significantly preserving model safety, offering an easy and robust solution for fine-tuning safety-aligned LLMs. Our key contributions are as follows: Our key contributions are as follows:
  • We propose a simple and effective method that improves downstream task performance while reducing the Attack Success Rate (ASR) by up to 30%.
  • We conduct a comprehensive evaluation using three models, four downstream tasks, four merging methods, and two safety benchmarks, demonstrating the robustness of our method in preserving model safety.
@@ 2 Related Work
2.1 Jailbreaking LLMs through Fine-Tuning
Despite significant efforts to ensure the safety of Large Language Models (LLMs) Ouyang et al. (2022); Rafailov et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024), adversarial attacks known as "jailbreaking" remain a major threat Xu et al. (2024); Yi et al. (2024). Additionally, studies Yang et al. (2023); Qi et al. (2024); Zhan et al. (2024) have shown that fine-tuning safety-aligned LLMs can degrade their safety, enabling these models to generate harmful responses.
A common mitigation strategy involves incorporating additional safety data during fine-tuning Qi et al. (2024); Bianchi et al. (2024). Other approaches use complex training or post-training techniques Huang et al. (2024b); Zong et al. (2024); Huang et al. (2024c). However, these methods are often limited by data availability and high computational costs. To overcome these challenges, we propose a simple yet effective method that mitigates safety degradation by merging the base model with the fine-tuned model, without requiring additional safety data or significant computational resources.
@@ 2.2 Model Merging
Model merging combines multiple models into a single unified model. A straightforward method averages the weights of different models Wortsman et al. (2022a). Advanced techniques such as SLERP White (2017), DARE Yu et al. (2024), and Model Stock Jang et al. (2024) have been developed to enhance the merging process.
Another approach leverages task vectors Ilharco et al. (2023), enabling flexibility and composability through arithmetic operations for various applications Huang et al. (2024a); Su et al. (2024). Some studies use task vectors to improve LLM safety by fine-tuning additional safe or harmful models to extract a safety vector Bhardwaj et al. (2024); Hazra et al. (2024); Hsu et al. (2024) and determining merging ratios based on safety data Hammoud et al. (2024). In contrast, our method utilizes the inherent safety of the base model, significantly reducing the need for additional safety data.
@@ 3 Methodology
To maintain the safety of safety-aligned LLMs after fine-tuning without relying on additional safety data, we propose a two-step approach. First, we fine-tune the base model on a target downstream task. Second, we merge the base model with the fine-tuned model by interpolating their weights. This merging step effectively mitigates safety degradation introduced during fine-tuning by leveraging the original safety-aligned LLMs.
Step 1: Fine-Tuning the Large Language Model
We fine-tune the base model on downstream tasks. For each task $t$, given an instruction $x^{t}$ and the corresponding response $y^{t}$, we optimize the model using the following loss function:
$\mathcal{L}_{FT}=-\log f_{\theta}(y^{t}|x^{t})$
where $f$ represents the language model parameterized by $\theta$.
Step 2: Merging the fine-tuned LLM with the base model
After fine-tuning, we merge the base model parameters ($\theta_{base}$) with those of the fine-tuned model ($\theta_{t}$) using the interpolation formula:
$\theta_{merged}=(1-\lambda)\theta_{base}+\lambda\theta_{t}$
Here, $\theta_{merged}$ represents the parameters of the merged model, and $\lambda\in[0,1]$ determines the contribution of the fine-tuned model. The ratios are normalized to ensure they sum to 1. In our experiments, we explore various merging techniques based on Equation 3, including methods that adjust model parameters before merging and strategies to identify the optimal $\lambda$.
@@ 4 Experimental Setups
In this section, we describe the experimental setup to address the following research questions: Q1: Can merging fine-tuned models with their base models prevent safety degradation after fine-tuning on different downstream tasks? Q2: How do different merging methods perform? Q3: What is the impact of $\lambda$ on downstream task performance and safety? Q4: Does model merging work on multiple models?
4.1 Downstream Tasks
We evaluate our proposed method on four downstream tasks: reasoning, medical assistance, code generation, and tool usage proficiency.
@@ 4.1 Downstream Tasks
Reasoning is improved with Chain-of-Thought data from the Flan V2 dataset Longpre et al. (2023) and is evaluated on the Big Bench Hard (BBH) dataset Suzgun et al. (2023). Medical assistance uses patient-doctor interactions from the ChatDoctor dataset Li et al. (2023). Code generation uses the MagiCoder dataset Wei et al. (2024) and is assessed via the HumanEval dataset Chen et al. (2021). Tool usage proficiency uses the OpenFunctions dataset Patil et al. (2023) for enhancing API call generation. For medical assistance and tool usage proficiency, response similarity to reference answers is measured using BERTScore [^1] Embedding extraction from the 40th layer of microsoft/deberta-xlarge-mnli. Zhang* et al. (2020). See Appendix A for more details on the downstream tasks.
4.2 Safety Evaluation
Safety is assessed using harmful instructions from the AdvBench dataset Chen et al. (2022) and the HEx-PHI dataset Qi et al. (2024). AdvBench contains 520 harmful instructions and HEx-PHI comprises 330 harmful instructions. We employ the WildGuard safety classification model Han et al. (2024), which demonstrates performance comparable to GPT-4 OpenAI et al. (2024). The Attack Success Rate (ASR) is used as the metric for both datasets. Further details on safety evaluation are provided in Appendix B.
4.3 Large Language Models
Our experiments involve two LLM families: Llama-3 Grattafiori et al. (2024) and Gemma Team et al. (2024). We use their instruct-tuned versions, which are aligned with human preferences.
Within the Llama-3 family, we evaluate Llama-3-8B-Instruct and Llama-3.1-8B-Instruct. For the Gemma family, we use the Gemma-2B-it model. Each model is fine-tuned with three different seeds using LoRA Hu et al. (2022) for each downstream task. Additional details on training and inference setups are provided in Appendix C.
4.4 Merging Methods
We evaluate three merging methods: Linear Merging, where models are directly combined by interpolating their weights following Equation 3, Model Stock Jang et al. (2024), SLERP White (2017), and DARE Yu et al. (2024). For Linear Merging, SLERP, and DARE, each fine-tuned model is merged with the base model using $\lambda$ selected based on validation set performance on the downstream tasks. Specifically, in the Model Stock approach, three models fine-tuned with different seeds are merged with the base model. More details about the methods and implementations are provided in Appendix D.
5 Results and Discussion
In this section, we present the results addressing our research questions. From Q1 through Q4, we focus primarily on the Llama-3-8B-Instruct model, other models are discussed in the context of Q5.
Q1: Can merging fine-tuned models with their base models prevent safety degradation after fine-tuning on different downstream tasks?
[…]
Table 1: Results across four tasks. The table reports the average scores of three models, which are fine-tuned with different seeds, for FT and Linear Merging. Performance indicates task-specific effectiveness measured by the respective metrics.
The results for Llama-3-8B-Instruct are presented in Table 1. While the fine-tuned model (FT) demonstrates improved performance on the downstream task, it significantly compromises safety. However, merging the fine-tuned model with the base model substantially enhances safety across all downstream tasks, as highlighted in the green cells in Table 1 and subsequent tables. Although the merged model does not fully match the base model in safety, it substantially reduces the ASR on both the AdvBench and HEx-PHI safety datasets. Notably, on the medical assistance task, the ASR is reduced by over 30%, bringing safety closer to the base model’s level. This result underscores the base model’s stronger inherent safety features. Additionally, Linear Merging improves downstream task performance across all downstream tasks, consistent with the findings of Wortsman et al. (2022b). Given the pronounced effect on the medical assistance task, subsequent experiments for questions Q2 through Q4 primarily focus on this task.
Q2: How do different merging methods perform?
[…]
Table 2: Results of various merging methods on the medical assistance task. The results show that different merging methods are beneficial for the safety of the fine-tuned model while also enhancing task performance.
The performance of alternative merging methods is shown in Table 2. Both SLERP and DARE effectively reduce ASR on AdvBench and HEx-PHI. The results emphasize the strong performance of Linear Merging, demonstrating that simpler approaches can effectively mitigate safety degradation. This finding suggests that Linear Merging can be a viable and computationally efficient alternative for practical applications.
Q3: What is the impact of $\lambda$ on downstream task performance and safety?
Figure 2: Impact of $\lambda$ on downstream task performance and safety. These plots illustrate how varying $\lambda$ affects the performance of the Medical Assistance task (left y-axis) and the ASR for the HEx-PHI dataset (right y-axis) across three merging methods. Results for AdvBench are shown in Figure 3 in the Appendix.
[…]
Table 3: Results of Llama-3.1-8B-Instruct and Gemma-2B-it The results show that our method can generalize to more models, restoring safety while maintaining task performance.
Figure 2 illustrates the impact of $\lambda$ on downstream task performance and model safety. As $\lambda$ increases, downstream task performance improves, but the ASR also rises, indicating a trade-off between performance and safety. However, the ASR remains lower than that of the SFT model. The optimal $\lambda$ is observed around $0.5\sim 0.6$, suggesting that more evenly combining the weights of the two models yields better downstream task performance while maintaining safety. Additionally, the rate of ASR increase for Linear Merging, as shown in Figures 2 and 3, is slower than that of SLERP and DARE, further demonstrating that Linear Merging is a more practical method in real-world applications.
Q4: Does model merging work on multiple models?
To evaluate the generalizability of model merging, we test Llama-3.1-8B-Instruct and Gemma-2B-it, with results shown in the Table 3. For these models, safety degradation on both benchmarks after fine-tuning is relatively mild. Nevertheless, model merging effectively restores safety without significantly compromising downstream task performance, demonstrating the method’s applicability across different LLMs.
@@ 6 Conclusion
We propose a two-step method to address safety degradation in aligned LLMs by merging the weights of pre- and post-fine-tuned models. This approach allows aligned LLMs to acquire new downstream task capabilities while preserving their original safety features, without the need for additional safety data. Experiments demonstrate its effectiveness across diverse downstream tasks, models, and merging techniques, providing a simple yet robust solution for adapting aligned LLMs.
@@ 7 Limitations
In our experiments, we evaluate only on benign data from the reasoning, medical assistance, code generation, and tool-using proficiency tasks, leaving out other areas such as law, finance, and others. While Section 5 examines the efficacy of our methods on these four downstream tasks, the performance of aligned models fine-tuned on other domains, languages, or datasets contaminated with harmful examples remains an area for future investigation. Furthermore, our methods are tested only on models with sizes of 2B and 8B from two model families. The efficacy of the approach on larger models and different model families is uncertain and warrants further exploration.
@@
Due to the higher cost of methods like LLM-as-Judge Chiang and Lee (2023); Liu et al. (2023), which involve using LLMs such as GPT-4 OpenAI et al. (2024) to evaluate the safety of model responses, we utilize an alternative approach by employing a classifier, WildGuard Han et al. (2024), to classify model responses as safe or unsafe. While this reduces costs, WildGuard has limitations: it struggles with more complex instructions, may produce false positives or false negatives, and provides less detailed evaluations. Consequently, we are unable to analyze which types of harmful instructions the models are more vulnerable to or which types are most effectively defended against after applying our method. More detailed safety analysis is left for future work.

### v2

@@ Abstract
Fine-tuning large language models (LLMs) for downstream tasks often leads to catastrophic forgetting, notably degrading the safety of originally aligned models. While some existing methods attempt to restore safety by incorporating additional safety data, the quality of such data typically falls short of that used in the original alignment process. Moreover, these high-quality safety data are generally inaccessible, making it difficult to fully recover the model’s original safety. We ask: How can we preserve safety while improving downstream task performance without additional safety data? We show that simply merging the weights of pre- and post-fine-tuned models effectively mitigates safety degradation while enhancing performance. Experiments across different downstream tasks and models validate the method’s practicality and effectiveness.
@@ 1 Introduction
The rapid advancement and increasing accessibility of Large Language Models (LLMs) necessitate a critical focus on aligning these technologies with human values, cultural norms, and trustworthiness Huang et al. (2023). To address these challenges, researchers and developers have introduced safety techniques such as preference tuning Ouyang et al. (2022); Rafailov et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024), aimed at preventing LLMs from generating harmful or inappropriate content. Many applications now leverage safety-aligned models as foundation models—referred to as aligned models in this paper—to further customize for downstream tasks via supervised fine-tuning (SFT) Chung et al. (2024).
Figure 1: Beyond standard SFT for downstream task adaptation, we can effectively mitigates safety degradation by combining the aligned and the fine-tuned model.
However, recent studies Yang et al. (2023); Qi et al. (2024); Zhan et al. (2024) highlight a critical challenge: fine-tuning aligned models can degrade their safety, even when using benign datasets. To address this issue, mainstream approaches often incorporate additional safety data during fine-tuning Qi et al. (2024); Bianchi et al. (2024). However, since the original safety data used to align LLMs are rarely available, surrogate data are typically generated by other LLMs-raising concerns about quality, and the potential for alignment drift.
In this paper, we demonstrate a simple yet effective method for improving downstream task performance while mitigating safety degradation. As illustrated in Figure 1, our approach consists of two steps: (1) fine-tune the aligned model on the downstream task, and (2) merge the aligned model with the fine-tuned model. We evaluate this strategy across various models and downstream tasks. Experimental results show that this method consistently enhances downstream task performance while substantially preserving model safety, offering a simple and robust solution for fine-tuning safety-aligned LLMs. Our key contributions are:
  • We show that a simple merging strategy can improve downstream task performance while lowering the Attack Success Rate (ASR).
  • We conduct extensive evaluations across three LLMs, four downstream tasks, and two safety benchmarks, demonstrating the robustness of our method in preserving model safety.
@@ 2 Related Work
2.1 Catastrophic Forgetting and Safety Degradation in LLMs
LLMs are commonly aligned with human preferences to ensure safety and reduce the likelihood of generating harmful content Ouyang et al. (2022); Rafailov et al. (2023); Grattafiori et al. (2024); OpenAI et al. (2024). However, recent studies have shown that this safety alignment can be significantly compromised after fine-tuning on downstream tasks Yang et al. (2023); Qi et al. (2024); Zhan et al. (2024). This degradation is often attributed to catastrophic forgetting Kirkpatrick et al. (2017); Li and Lee (2024); Luo et al. (2025), a well-known challenge in post-training scenarios where a model forgets previously acquired knowledge when adapting to new tasks.
To mitigate this issue, prior work has proposed several approaches. One line of work augments fine-tuning with additional safety data Qi et al. (2024); Bianchi et al. (2024); Zong et al. (2024), aiming to reinforce desirable behaviors through curated examples. Another line of work leverages self-distillation, where the model generates synthetic training data, and fine-tuning on this data has been shown to reduce harmful tendencies Yang et al. (2024). In addition, some studies explore incorporating regularization strategies during training, often combined with additional safety data, to constrain harmful deviations Huang et al. (2024c, d). Others adopt post-hoc re-alignment methods, such as Huang et al. (2024b) and Yi et al. (2024), which utilize additional safety data to identify safety-related masks and subsequently apply these masks in their re-alignment processes.
However, these methods either require synthesizing large amounts of safety data or incur significant computational overhead. In contrast, the approach proposed in this paper avoids both additional data requirements and extra training costs, offering a more efficient alternative for maintaining safety.
@@ 2.2 Model Merging
Model merging combines multiple models into a single unified model. A straightforward approach is to average the weights of different models Wortsman et al. (2022a), while variant techniques include SLERP White (2017) and DARE Yu et al. (2024).
Another line of work explores task vectors Ilharco et al. (2023), typically computed as the difference between a fine-tuned model and its base. These vectors enable composable transformations across tasks Huang et al. (2024a); Su et al. (2024) and have been extended to construct “safety vectors” from separate safe or harmful models to prevent safety degradation. Bhardwaj et al. (2024), Hazra et al. (2024), and Wu et al. (2025) adopt a similar approach: they first fine-tune an aligned model on harmful data to obtain a harmful variant, then compute a safety vector as the parameter difference between the aligned and harmful models, capturing the directional shift introduced by safety alignment. In contrast, Hsu et al. (2024) avoids additional fine-tuning but assumes access to a pre-alignment checkpoint to derive the safety vector for subsequent alignment, which is not always publicly available. In contrast, our method uses only aligned models and fine-tuned models, making it widely applicable, and demonstrates that safety can be restored without extra safety data.
The proposed approach is similar to WiSE-FT Wortsman et al. (2022b), which also interpolates between the base model and its fine-tuned variant. However, WiSE-FT is applied to computer vision, not LLMs, and is not aimed at preserving safety.
@@ 3 Methodology
Our method comprises just two stages: (1) fine-tuning the aligned model on a target downstream task, and (2) merging the original aligned model with the fine-tuned model by interpolating their weights. Despite its simplicity, this merging strategy effectively mitigates the degradation in safety commonly observed following fine-tuning, while preserving performance on the target downstream task, without requiring additional data.
Step 1: Supervised Fine-Tuning of the Large Language Model
We fine-tune the aligned model with parameters $\theta_{\text{base}}$ on a given task $t$, resulting in a task-specific model $\theta_{t}$. For each task $t$, given an instruction $x^{t}$ and its corresponding response $y^{t}$, we minimize the negative log-likelihood:
$\mathcal{L}_{FT}=-\log f_{\theta}(y^{t}\mid x^{t})$
where $f_{\theta}$ denotes the language model parameterized by $\theta$.
Step 2: Merging the Fine-Tuned Model with the Aligned Model
After fine-tuning, we merge the parameters of the aligned model ($\theta_{\text{base}}$) with those of the fine-tuned model ($\theta_{t}$) via linear interpolation:
$\theta_{\text{merged}}=(1-\lambda)\theta_{\text{base}}+\lambda\theta_{t}$
Here, $\theta_{\text{merged}}$ denotes the parameters of the merged model, and $\lambda\in[0,1]$ controls the relative contribution of the fine-tuned model. Eq. 3 is the formulation for the native linear merging method; other advanced merging methods can also be applied.
@@ 4 Experimental Setups
Downstream Tasks
@@
We conduct experiments on four downstream tasks: reasoning, medical assistance, code generation, and tool usage proficiency. Reasoning is enhanced using Chain-of-Thought data from the Flan V2 dataset Longpre et al. (2023) and evaluated on the Big Bench Hard (BBH) dataset Suzgun et al. (2023). Medical assistance uses patient-doctor dialogues from the ChatDoctor dataset Li et al. (2023). Code generation is trained on the MagiCoder dataset Wei et al. (2024) and evaluated using the HumanEval benchmark Chen et al. (2021). Tool usage proficiency leverages the OpenFunctions dataset Patil et al. (2023) to improve API call generation. For medical assistance and tool usage proficiency, response similarity to reference answers is measured using BERTScore [^1] Embeddings extracted from the 40th layer of microsoft/deberta-xlarge-mnli. Zhang* et al. (2020). See Appendix A for additional details on the downstream tasks.
Safety Evaluation
We assess safety using harmful instructions from the AdvBench Chen et al. (2022) and HEx-PHI Qi et al. (2024) datasets. Following prior works that use safety classifiers to automatically detect harmful content Xie et al. (2025); O’Brien et al. (2024), we adopt WildGuard Han et al. (2024), a classifier shown to perform comparably to GPT-4 OpenAI et al. (2024). We report the Attack Success Rate (ASR) as the primary evaluation metric. Details of the evaluation setup are provided in Appendix B.
Large Language Models
Our experiments involve several LLMs, including LLaMA-3-8B-Instruct Grattafiori et al. (2024), Gemma-2-2B-It Team et al. (2024), and Qwen2.5-7B-Instruct Team (2024), along with additional model sizes when noted. We use the instruct-tuned variants of all models, which are aligned with human preferences. Each model is fine-tuned on each downstream task using LoRA Hu et al. (2022) with three different random seeds. The reported downstream task performance and ASR are averaged across these three runs. Additional details of experiment are provided in Appendix C.
Baselines
Unlike most existing methods aimed at mitigating safety degradation in LLMs after fine-tuning, our proposed approach requires neither additional data nor further training. Given the absence of comparable safety alignment techniques, we evaluate our method’s efficacy in preserving the safety attributes of the originally aligned model post fine-tuning by benchmarking it against two prevalent regularization techniques: Dropout Srivastava et al. (2014) and Weight Decay Loshchilov and Hutter (2019). Similar to our approach, these regularization methods do not necessitate extra data or further training. The hyperparameters for these techniques are selected based on validation set performance on downstream tasks.
Merging Methods
In Section 5, we used Linear Merging, which combines models via direct interpolation as defined in Eq. 3, as the merging method. Two advanced merging methods—SLERP and DARE—are also applied. Their results are provided in Appendix E. For all methods, we merge each fine-tuned model with the aligned model using an interpolation factor $\lambda$ selected based on validation set performance.
5 Results
5.1 Can model merging mitigate safety degradation after fine-tuning?
Figure 2: Pareto analysis of downstream task performance and ASR on AdvBench across different models and tasks. Each dot represents a model configuration, with different hyperparameter settings (weight decay coefficient, dropout rate, or merging interpolation coefficient) for the same method shown in the same color. For clarity, we connect the dots of our method in ascending order of their coefficients. Dots with dark edges indicate the best-performing models on the validation set for each method.
Figure 2 presents a Pareto analysis of task performance and ASR on AdvBench across different models and tasks. We observe that SFT consistently leads to safety degradation, with higher ASR across all settings compared to the original aligned model. While Dropout and Weight Decay offer slight improvements in ASR, they are generally insufficient to restore the safety of the aligned model.
In contrast, the proposed approach consistently achieves a better balance between performance and safety. It often reduces ASR to levels near that of the aligned model while maintaining—or even improving—task performance. The smooth Pareto fronts formed by merging indicate controllable trade-offs, making it an effective solution for mitigating safety loss after fine-tuning. Figure 5 shows the HEx-PHI results, and results of different merging methods can be found in Table 1 in Appendix E.
5.2 How does model merging perform across different model sizes?
Figure 3: Performance and ASR change across model sizes. This figure shows results for Qwen2.5 at 1.5B, 3B, and 7B (top), and Gemma-2 at 2B and 9B (bottom).
Figure 4: Accuracy of LLaMA-3 on IFEval. This figure shows results of LLaMA-3-8B-Instruct fine-tuned on downstream tasks on IFEval. Fine-tuning reduces instruction-following ability, especially for Reasoning and Medical tasks. Merging with the Aligned model helps recover this ability close to the original level.
Luo et al. (2025) noted that larger models may suffer more from catastrophic forgetting. We extend this analysis to safety degradation and evaluate how model merging performs across different model sizes. Figure 3 shows the average changes in performance and ASR across all downstream tasks for the Qwen2.5 and Gemma-2 model families, comparing SFT and the proposed approach against their respective aligned models. Both methods improve task performance, with larger models generally achieving greater gains. However, safety degradation shows no consistent trend: smaller Qwen models degrade more, while larger Gemma models are more affected. This suggests that safety degradation is not solely determined by model size. Nonetheless, the proposed approach consistently mitigates safety degradation across different scales.
5.3 Can model merging help preserve other capabilities of the aligned model?
While our method mitigates safety degradation, we also investigate whether it preserves other capabilities of the aligned model that are lost due to catastrophic forgetting. Since we fine-tune instruct-tuned variants, we evaluate whether instruction-following ability is retained. Figure 4 shows the performance of LLaMA-3-8B-Instruct on the instruction-following benchmark IFEval Zhou et al. (2023). Both prompt and instruction accuracy decline after fine-tuning, with the largest drops observed in the reasoning and medical tasks. The proposed approach substantially restores performance to the level of the aligned model, indicating that merging can also preserve instruction adherence.
@@ 6 Conclusion
We present a simple yet effective method to address the safety degradation that often occurs when adapting LLMs to downstream tasks, without requiring additional safety data or auxiliary models. The method also preserves capabilities such as instruction-following, making it a practical and scalable solution for adapting LLMs to new tasks.
@@ 7 Limitations
In our experiments, we evaluate only on benign data from four task domains: reasoning, medical assistance, code generation, and tool-using proficiency. Other important areas such as law, finance, or multilingual tasks remain unexplored. While Section 5 shows the effectiveness of our method on the selected downstream tasks, its generalizability to other domains, languages, or datasets that may contain harmful content remains an open question. Additionally, we evaluate models with sizes ranging from 1.5B to 9B across three model families. The effectiveness of our approach on larger models or different model architectures warrants further investigation.
@@
Due to the high computational and financial cost of human-aligned safety evaluation methods such as LLM-as-Judge Chiang and Lee (2023); Liu et al. (2023), which require using large proprietary models like GPT-4 OpenAI et al. (2024), we instead adopt WildGuard Han et al. (2024), a lightweight open-source safety classifier. WildGuard is shown to perform competitively with GPT-4 on multiple safety detection tasks and offers a reproducible, low-cost alternative suitable for large-scale evaluations.
However, this classifier-based approach has several limitations. First, WildGuard may struggle with complex or subtle harmful instructions, potentially leading to both false positives and false negatives. Second, it provides only binary or coarse-grained outputs (e.g., “harmful” or “safe”), without offering finer distinctions such as the category of harm, the severity of the risk, or whether the model’s refusal was appropriate or evasive.
Consequently, while WildGuard enables efficient and scalable evaluation, it constrains the depth of our safety analysis. Future work could incorporate more fine-grained multi-label safety classifiers, adversarial evaluation pipelines, or hybrid setups involving human or LLM-as-Judge verification to better capture the nuanced impact of model merging on safety behavior.
Jailbreak Attacks
Our work focuses on safety degradation that arises from fine-tuning aligned LLMs on benign tasks, which we consider a case of catastrophic forgetting. As such, we evaluate whether models produce harmful outputs when directly prompted with harmful instructions, rather than testing resistance to specific jailbreak strategies. We do not include jailbreak-style attacks Xu et al. (2024) in our evaluation due to two reasons: (1) Our primary goal is to study alignment loss under standard fine-tuning, not adversarial robustness; and (2) jailbreak evaluations typically require separate prompting strategies and adversarial instruction crafting pipelines, which are beyond the scope of this study. Future work can extend our framework to examine the impact of merging on robustness against jailbreak attacks.